OpenStudy (anonymous):

can anybody show me step-by-step on how to find the value of the unknown variables? please.. problem: find value of x, y, and z. x+y=24 y+z=26 x+z=28 ASAP. .

OpenStudy (anonymous):

put the @ before so he can see it in his notification... I'm trying it out I'll get back to you if I get it

OpenStudy (anonymous):

@ishaan94 help. .

OpenStudy (anonymous):

ok, I managed to get an answer, but I feel that there should be a much more direct answer, but this is what I got.:

OpenStudy (anonymous):

can you please show me?

OpenStudy (anonymous):

From algebra we know we can tweak with equations... so: the 3 normal equations x+y=24 y+z=26 x+z=28 we get x=24-y x=28-z y=24-x y=26-z z=26-y z=28-x now, to find out a little more information, I played with it again. since x=24-y and 28-z then, 24-y=28-z so then this gives when we move to integers to combine like terms in one side; -4-y=z or 4+y=z which means that z is 4 more units bigger than whatever y is... now, y=26-z = 24-x 26-z=24-x z=2+x z is 2 more than whatever x is and finally, z=26-y=28-x 26-y=28-x 2+y=x so x is 2 units bigger than whatever y is so... once again the info x is 2 more than y z is 2 more than x z is 4 more than y from there we just see for first equation y+x=24 and from the information obtained we get y=11 and x= 13 and now that you know these, getting z is a breeze

OpenStudy (anonymous):

how about the value of z?

OpenStudy (anonymous):

well since we know that z is 2 more than x and 4 more than y all you have to do is add 2 to x or 4 to y to get z... which is 15 11+4=15 13+2=15 so z=15

OpenStudy (anonymous):

and you can double check by plugging it into the formulas

OpenStudy (anonymous):

thanks. . .

OpenStudy (anonymous):

yea no problem, I do believe this was the primeval methods used to solve this equation lol.. but I don't know any modern ones... except using matlab, which is a software you might learn to use if you take engineering courses... if you find out a shorter way to do it.. let me know

OpenStudy (anonymous):

lol. . the question is, if i can? lol. . but' anyway thanks again!